javonarnold:

Ruiz needs to buy a countertop for a kitchen. He calculated the area to be 16 square feet. The actual area is 15.4 square feet. What is Ruiz's percentage error?

jhonyy9:

so 15,4 ft^2 mean the 100% area 16 -15,4 = 0,6 and now just we use the rule of 3 simple - write in this way 15,4 ft^2 mean 100% 0,6 ft^2 mean x% cross multiplie and calcule the x value x = 0,6 *100/15,4 = 3,89% so = aprox. 4%

xxDeppressionxx:

16 - 15.4 = 0.6, 0.6 / 15.4 = 0.038961039, 0.038961039 * 100 = 3.8961039 so 3.9% is your answer which is almost = to the whole number 4